# A Systematic Search for Spectral Hardening in Blazar Flares with the Fermi-Large Area Telescope

### v1

@@ Abstract
Context. Blazars, a subclass of active galactic nuclei (AGN), are among the most powerful and variable -ray sources in the universe. They emit non-thermal radiation across the electromagnetic spectrum in the form of relativistic jets, characterized by rapid flux and polarization variability. Extreme high synchrotron-peaked blazars (EHSPs), with synchrotron peaks exceeding Hz, are crucial for understanding the full range of blazar phenomena and testing models of jet physics. However, the number of known extreme blazars is small, and, as a result, this class of objects remains poorly studied.
@@
Methods. We analyze a sample of 138 blazars selected from the 4FGL-DR2 catalog with high synchrotron peak frequencies and well-sampled light curves. Flaring periods are selected using Bayesian Block analysis. Each flare is then analyzed through -ray spectral fitting with both power-law and broken power-law models to identify potential spectral hardening. The significance of spectral hardening is assessed using a test statistic, TShardening, based on the likelihood ratio of the two spectral models.
Results. We identify two flaring episodes with indications of spectral hardening, one in 4FGL J0238.43116 and another in PKS 2155304, the latter detected independently by both selection methods but referring to the same flaring period. This number of candidate events is consistent with expectations from statistical fluctuations, suggesting that spectral hardening is, at most, a rare occurrence in -ray blazars. These results provide the first population-level constraint on the frequency of such events (¡ 0.1%). The scarcity of events reinforces the notion that the dominant blazar emission mechanism is well described by smoothly varying power-law spectra across the Fermi-LAT range, with sharp spectral hardenings representing rare deviations likely tied to exceptional jet conditions or transient physical processes. Although these flares show notable spectral changes, their statistical significance remains modest and motivate future multi-wavelength studies to assess whether these rare flares reflect genuinely distinct physical processes within blazar jets.
@@
Blazars, a subclass of active galactic nuclei (AGN) with relativistic jets aligned close to our line of sight, are among the most powerful and variable sources of -rays in the universe (e.g., Urry & Padovani 1995). These jets emit non-thermal radiation across the entire electromagnetic spectrum, from radio to -rays, driven by relativistic particles within the jet (e.g., Ulrich et al. 1997). Blazars are characterized by rapid flux variability, high polarization, and apparent superluminal motion (e.g., Jorstad et al. 2001; Marscher 2008).
@@
Blazars are classified into BL Lac objects (BL Lacs) and flat-spectrum radio quasars (FSRQs) based on their emission line properties, with weak or absent emission lines in the former and strong lines in the latter in their optical spectra (e.g., Urry & Padovani 1995; Padovani et al. 2017). There is also an additional sub-classification based on the synchrotron peak frequency (e.g., Abdo et al. 2010b). Extreme high synchrotron-peaked blazars (EHSPs) are particularly noteworthy, with synchrotron peak frequencies exceeding Hz (e.g., Costamante et al. 2001; Paliya et al. 2019). These extreme blazars are critical for understanding the full range of blazar phenomena and for testing models of jet physics, particularly in the context of the blazar sequence, which attempts to order blazars by luminosity and synchrotron peak frequency (e.g., Fossati et al. 1998; Ghisellini & Tavecchio 2008; Ghisellini et al. 2017). However, the blazar sequence has been questioned, with suggestions that observational biases may influence the perceived trends (e.g., Giommi et al. 2012; Chang et al. 2019; Keenan et al. 2021). The blazar sequence framework is helpful for understanding the diverse properties of blazars and how they may be related to the orientation and physical conditions within their relativistic jets. However, the most extreme blazars have not been extensively studied, and their properties remain poorly categorized. Observing these sources is essential for improving our understanding of the blazar sequence and blazar evolution. While the exact number is challenging to determine due to variations across surveys, only a limited number of these sources are known (e.g., Chang et al. 2019; Nievas Rosillo et al. 2022). Note that a blazar may exhibit the properties of an EHSP temporarily during flaring states, as these events can shift the synchrotron peak toward higher energies. Therefore, identifying such transient states is crucial for a complete characterization of EHSPs (Valverde et al. 2020; Abe et al. 2024).
@@
Furthermore, according to Biteau et al. (2020) and Prandini & Ghisellini (2022), a majority of known TeV blazars show extreme synchrotron peaks. Therefore, they conclude that TeV blazars may have a greater tendency to display extreme behaviors in terms of SED peaks. As very high energy (VHE, GeV) extragalactic particle accelerators, extreme blazars represent prime candidates for exploring multi-messenger correlations with neutrinos and cosmic rays (e.g., Costamante et al. 2001). Given their hard spectrum at higher energies, these blazars are important for studying cosmic -ray photon propagation, including the extragalactic background light and cosmological research (e.g., Domínguez et al. 2013; Domínguez & Ajello 2015; Domínguez et al. 2019; Saldana-Lopez et al. 2021; Domínguez et al. 2024). They may also be the dominant component of the extragalactic TeV background (e.g., Giommi & Padovani 2015; Ajello et al. 2015). All this makes understanding EHSPs a fundamental topic and desirable targets for current and future VHE telescopes such as the Cherenkov Telescope Array Observatory (CTAO; Acharya et al. 2017). In summary, a complete characterization of extreme blazars is currently lacking, and the discovery of these blazars holds significant importance for the field of VHE astrophysics.
@@
In this study, we aim to identify and characterize the most extreme -ray blazars using data from the Large Area Telescope (LAT) on board the Fermi Gamma-ray Space Telescope. Our focus is on detecting spectral hardening, a feature where the -ray spectrum becomes harder at higher energies, particularly during flaring episodes. Such hardening could indicate a transition between the synchrotron and IC peaks, or other extreme physical conditions within the jet. Previous studies have reported similar features, but their occurrence is rare and often linked to specific flaring events (Abdo et al. 2010b; Zeng et al. 2022). By systematically searching for this feature in a large sample of blazars, we aim to enhance our understanding of the most extreme blazar behaviors and their underlying physical mechanisms.
@@ 2.1 Source Selection
In this section, we describe the criteria used for selecting our sample of blazars. The sources are drawn from the 4FGL-DR2 catalog (Ballet et al. 2020), for which light curves (LCs) have been previously computed and utilized by Peñil et al. (2022), Peñil et al. (2024b), Peñil et al. (2024a), Rico et al. (2025) and Peñil et al. (2025). These LCs include all AGN types (3308 AGNs) and cover the first 12 years of Fermi-LAT data in 28-day binning. We apply the following selection criteria: (1) High-latitude Fermi-LAT sources , to minimize contamination from diffuse emission in the Galactic plane. (2) A synchrotron peak frequency greater than Hz, based on values from the Fourth Catalog of Active Galactic Nuclei detected by the LAT (4LAC) - DR3 catalog (Ajello et al. 2022) or Yang et al. (2022), with preference given to the latter reference. The reason is that Yang et al. (2022) developed an analysis specially focused on synchrotron peaks. Using these two criteria, we obtain 365 sources. Finally, we require that the LCs contain less than 50% upper-limit (UL) bins, meaning that more than half of the time bins must be detections111We define a time bin as a detection if its Test Statistic (TS) exceeds 1, and otherwise consider it an UL. See Sec.3.1 for further discussion. to allow for reliable flare identification (e.g., Peñil et al. 2020). This leaves a final sample of 138 blazars.
The selection of sources with synchrotron peak frequencies above Hz is driven by our aim to identify blazars exhibiting spectral hardening in the -ray band. Blazars with higher synchrotron peak frequencies are more likely to show spectral breaks within the energy range detectable by Fermi-LAT, as their synchrotron emission extends further into higher energies, increasing the probability of encountering a spectral transition in the framework of synchrotron/SSC models (Ghisellini et al. 2017). Additionally, by selecting blazars with already high synchrotron peak frequencies, we reduce the need for large shifts during flaring states to observe these spectral breaks within the Fermi-LAT energy range.
@@ 2.2 Gamma-ray Data Reduction and Spectral Fitting
We perform the analysis of the -ray data using the Fermipy software package (version 1.3.1), an open-source Python package designed to analyze Fermi-LAT data (Wood et al. 2017). The analysis follows the criteria for a point source analysis, utilizing the latest Galactic and isotropic diffuse models, gll_iem_v07.fits and iso_P8R3_SOURCE_V3_v1.txt, respectively222https://Fermi.gsfc.nasa.gov/ssc/data/access/lat/BackgroundModels.html. The analysis is done using the latest version of the Fermi Science Tools (V2.2.0) and the instrument response function used is P8R3 SOURCE V3 (Atwood et al. 2013).
We conduct the spectral fitting over flaring periods, which are identified as described in Section 3.1. For each source in our sample, we define a region of interest (ROI) with a radius of 15 degrees centered on the source. The filter expression used is ‘(DATA QUAL¿0)(LAT CONFIG == 1)(angsep(RA source, DEC source, RA_SUN, DEC_SUN)¿15)’, where RA source and DEC source are the right ascension and declination of the source, and RA_SUN and DEC_SUN are those of the Sun. This selection ensures that only good time intervals (GTIs) are considered, and time periods when the source is within 15 degrees of the Sun are excluded to prevent contamination from solar emission.
The -ray sky model is constructed using the 4FGL-DR3 catalog, and the energy range considered is 50 MeV to 300 GeV. We use different PSF event types and zenith angle cut for different energy ranges to optimize angular resolution and sensitivity:
@@
This selection maximizes sensitivity to point sources like blazars, which is crucial for accurate spectral and spatial analyses.
@@
The normalization of the Galactic and isotropic diffuse component is left free. Additionally, the normalization parameter of catalog sources with a test statistic (TS) value greater than 25 within 9 degrees of the ROI center is set free. For sources with TS 500 within 12 degrees, both the normalization and spectral index are allowed to vary, accounting for long-term averaged source properties over the first 12 years of Fermi-LAT data. After an initial fitting, a source-finding step is implemented using gta.find_sources from Fermipy to identify additional unresolved sources in the ROI. Following this, the same TS-based criteria are reapplied to both catalog and newly identified sources during the analysis period, refining the model to account for temporal variability or new source activity. Finally, a global fit is performed to obtain the optimized model parameters.
@@
We then test the broken power-law (BPL) model as a candidate spectral function. Initially, we fit a power-law (PL) spectral function and record the likelihood of the fit, . Subsequently, we change the spectral model to a broken power-law (BPL) function, defined as:
@@
where is the break energy (in MeV), is the photon index for , and is the photon index for . We iterate the fit by fixing the break energy, , at 100 logarithmic intervals between 0.1 GeV and 10 GeV, calculating the likelihood, , at each step. The photon indices ( and ) are allowed to vary between -5 and 5. At each iteration, we calculate the value of TShardening = 2 ( - ), which quantifies the test statistics of spectral hardening from comparing the BPL model likelihood with the PL model likelihood. Note that this is similar, but more specific, to the TScurv used in the 4FGL catalog for determining whether a curved spectral function is preferred over a PL function (Abdollahi et al. 2020). And while creating the SEDs, we left free the parameters of all sources within 3 degrees. We finally select the that maximizes the TShardening profile.
@@
The choice of a BPL function is both statistically and physically motivated. Flaring periods are characterized by rapid spectral changes and generally short integration times, making smoother functions such as a logparabola less effective for constraining spectral hardening (e.g., Abdo et al. 2010a; Ackermann et al. 2011). Furthermore, a BPL, with an abrupt spectral break and two independent photon indices, better captures sudden changes in particle acceleration or cooling while providing a clearer quantification of spectral hardening (e.g., Tavecchio et al. 2009).

### v2

@@ Abstract
Context. Blazars, a subclass of active galactic nuclei (AGNs), are among the most powerful and variable -ray sources in the universe. They emit non-thermal radiation across the electromagnetic spectrum in the form of relativistic jets, characterized by rapid flux and polarization variability. High synchrotron-peaked blazars (HSPs) and extreme high synchrotron-peaked blazars (EHSPs), with synchrotron peaks exceeding Hz and Hz, respectively, are crucial for understanding the full range of blazar phenomena and testing models of jet physics. Yet, their understanding remains challenging.
@@
Methods. We analyzed a sample of 138 blazars selected from the 4FGL-DR2 catalog with high synchrotron peak frequencies and well-sampled light curves. Flaring periods were selected using Bayesian Block analysis. Each flare was then analyzed through -ray spectral fitting with both power-law and broken power-law models to identify potential spectral hardening. The significance of spectral hardening was assessed using a test statistic, TShardening, based on the likelihood ratio of the two spectral models.
Results. We identified two flaring episodes with indications of spectral hardening, one in 4FGL J0238.43116 and another in PKS 2155304, the latter detected independently by both selection methods but referring to the same flaring period. This number of candidate events is consistent with expectations from statistical fluctuations, suggesting that spectral hardening is, at most, a rare occurrence in -ray blazars. These results provide the first population-level constraint on the frequency of such events (¡ 0.1%). The scarcity of events reinforces the notion that the dominant blazar emission mechanism is well described by smoothly varying power-law spectra across the Fermi-LAT range, with sharp spectral hardenings representing rare deviations likely tied to exceptional jet conditions or transient physical processes. Although these flares show notable spectral changes, their statistical significance remains modest and motivates future multi-wavelength studies to assess whether these rare flares reflect genuinely distinct physical processes within blazar jets.
@@
Blazars, a subclass of active galactic nuclei (AGNs) with relativistic jets aligned close to our line of sight, are among the most powerful and variable sources of -rays in the universe (e.g., Urry & Padovani 1995). These jets emit non-thermal radiation across the entire electromagnetic spectrum, from radio to -rays, driven by relativistic particles within the jet (e.g., Ulrich et al. 1997). Blazars are characterized by rapid flux variability, high polarization, and apparent superluminal motion (e.g., Jorstad et al. 2001; Marscher 2008).
@@
Blazars are classified into BL Lac objects (BL Lacs) and flat-spectrum radio quasars (FSRQs) based on their emission line properties, with weak or absent emission lines in the former and strong lines in the latter in their optical spectra (e.g., Urry & Padovani 1995; Padovani et al. 2017). There is also an additional sub-classification based on the synchrotron peak frequency (e.g., Abdo et al. 2010). High synchrotron-peaked blazars (HSPs) and extreme high synchrotron-peaked blazars (EHSPs) are particularly noteworthy, with synchrotron peak frequencies exceeding Hz and Hz, respectively (e.g., Costamante et al. 2001; Paliya et al. 2019). These blazars are critical for understanding the full range of blazar phenomena and for testing models of jet physics, particularly in the context of the blazar sequence, which attempts to order blazars by luminosity and synchrotron peak frequency (e.g., Fossati et al. 1998; Ghisellini & Tavecchio 2008; Ghisellini et al. 2017). However, the blazar sequence has been questioned, with suggestions that observational biases may influence the perceived trends (e.g., Giommi et al. 2012; Chang et al. 2019; Keenan et al. 2021). The blazar sequence framework is helpful for understanding the diverse properties of blazars and how they may be related to the physical conditions within their relativistic jets. However, a comprehensive understanding of HSP and EHSP blazars remains challenging due to their intrinsic diversity and limited availability of sources with detailed and simultaneous multi-wavelength spectral coverage (e.g., Costamante et al. 2018; Foffano et al. 2019; Costamante 2020; Nievas Rosillo et al. 2022; MAGIC Collaboration et al. 2024; Sahu et al. 2025; Banerjee et al. 2025; Láinez et al. 2025). Observing these sources is essential for improving our understanding of the blazar sequence and blazar evolution. Note that a blazar may show the properties of an EHSP temporarily during flaring states, as these events can shift the synchrotron peak toward higher energies. Therefore, identifying such transient states is crucial for a complete characterization of EHSPs (Valverde et al. 2020; Abe et al. 2024).
@@
Furthermore, according to Biteau et al. (2020) and Prandini & Ghisellini (2022), a majority of known TeV blazars show extreme synchrotron peaks. Therefore, they conclude that TeV blazars may have a greater tendency to display extreme behaviors in terms of SED peaks. As very high energy (VHE, GeV) extragalactic particle accelerators, extreme blazars represent prime candidates for exploring multi-messenger correlations with neutrinos and cosmic rays (e.g., Costamante et al. 2001). Given their hard spectrum at higher energies, these blazars are important for studying cosmic -ray photon propagation, including the extragalactic background light and cosmological research (e.g., Domínguez et al. 2013; Domínguez & Ajello 2015; Domínguez et al. 2019; Saldana-Lopez et al. 2021; Domínguez et al. 2024). They may also be the dominant component of the extragalactic TeV background (e.g., Giommi & Padovani 2015; Ajello et al. 2015). All this makes understanding EHSPs a fundamental topic and desirable targets for current and future VHE telescopes such as the Cherenkov Telescope Array Observatory (CTAO; Acharya et al. 2017). In summary, despite numerous individual studies, a complete and uniform characterization of extreme blazars across a large sample remains limited, highlighting the value of systematic searches and population-level analyses (e.g., Biteau et al. 2020; Prandini & Ghisellini 2022).
@@
In this study, we used Fermi-LAT data to search for spectral hardening in a large sample of HSP blazars. Similar behavior has been reported in radio galaxies (Abdo et al. 2010; Brown et al. 2017; Rulten et al. 2020), in the HSP 1ES 0502+675 (Zeng et al. 2022), and in an FSRQ (Paliya et al. 2025), but it has not been systematically explored in HSP blazars, where synchrotron peaks may shift into the LAT band during flares. We employed Bayesian Block Analysis to identify robust flaring periods and test for the presence of spectral hardening by comparing fits of power-law versus broken power-law models to the gamma-ray spectra. Such spectral hardening could hint at complex conditions within blazar jets, such as transitions between different emission regions or electron populations, and deserves detailed multi-wavelength investigations (e.g., Abramowski et al. 2012; Dmytriiev et al. 2021). Previous studies have reported similar features, but their occurrence is rare and often linked to specific flaring events (Abdo et al. 2010; Zeng et al. 2022). By systematically searching for this feature in a large sample of blazars, we aim to enhance our understanding of the most extreme blazar behaviors and their underlying physical mechanisms.
@@ 2.1 Source selection
In this section, we describe the criteria used for selecting our sample of blazars. The sources were drawn from the 4FGL-DR2 catalog (Ballet et al. 2020), for which light curves (LCs) have been previously computed and utilized by Peñil et al. (2025), Peñil et al. (2024b), Peñil et al. (2024a), Rico et al. (2025), and Peñil et al. (2025). These LCs include all AGN types (3308 AGNs) and cover the first 12 years of Fermi-LAT data in 28-day binning. We applied the following selection criteria: (1) High-latitude Fermi-LAT sources , to minimize contamination from diffuse emission in the Galactic plane. (2) A synchrotron peak frequency greater than Hz, based on synchrotron peak frequencies taken from Yang et al. (2022), and from the 4LAC–DR3 catalog (Ajello et al. 2022) only when no value is available in the former. The reason is that Yang et al. (2022) developed an analysis specifically focused on synchrotron peaks, though with possible uncertainties due to non-simultaneous data, which also affect the 4LAC estimates (see references for details). Using these two criteria, we obtain 365 sources. Finally, we require that the LCs contain less than 50% upper-limit (UL) bins, meaning that more than half of the time bins must be detections111We define a time bin as a detection if its Test Statistic (TS) exceeds 1, and otherwise consider it an UL. See Sec.3.1 for further discussion. to allow for reliable flare identification (e.g., Peñil et al. 2020). We note explicitly that this condition of requiring fewer than 50% upper limits per flare biases our analysis toward brighter flares and more luminous sources. As a consequence, our results primarily characterize variability and spectral features in relatively bright events, and caution should be exercised when generalizing these findings to intrinsically faint HSP and EHSP sources. This results in a final sample of 138 blazars, with synchrotron peak frequencies taken from Yang et al. (2022) for 129 sources and from 4LAC-DR3 for 9 sources. Of these 138 blazars, 20 have synchrotron peak frequencies greater than Hz.
The selection of sources with synchrotron peak frequencies above Hz is motivated by our aim to identify blazars that may exhibit spectral hardening in the -ray band. Blazars with higher synchrotron peak frequencies are more likely to show spectral hardening within the energy range detectable by Fermi-LAT, as their synchrotron emission extends further into higher energies, increasing the probability of encountering a spectral transition in the framework of synchrotron/SSC models (Ghisellini et al. 2017). By selecting blazars with already high synchrotron peak frequencies, we also reduce the need for large shifts during flaring states to observe such spectral hardening within the Fermi-LAT band. This threshold ( Hz) includes both classical HSP and EHSP blazars. The rationale is to investigate whether spectral hardening observed during flaring episodes could occasionally correspond to synchrotron peak frequencies temporarily moving above Hz, thus producing transient EHSP-like states (e.g., Valverde et al. 2020; Abe et al. 2024). We emphasize that synchrotron peak values are used only for the initial source selection, while our -ray spectral analysis and identification of spectral hardening are carried out solely with Fermi-LAT data, independent of assumptions about synchrotron peak positions.
@@ 2.2 Gamma-ray data reduction and spectral fitting
We performed the analysis of the -ray data using the Fermipy software package (version 1.3.1), an open-source Python package designed to analyze Fermi-LAT data (Wood et al. 2017). The analysis followed the criteria for a point source analysis, utilizing the latest Galactic and isotropic diffuse models, gll_iem_v07.fits and iso_P8R3_SOURCE_V3_v1.txt, respectively222https://Fermi.gsfc.nasa.gov/ssc/data/access/lat/BackgroundModels.html. The analysis is done using the latest version of the Fermi Science Tools (V2.2.0) and the instrument response function used is P8R3 SOURCE V3 (Atwood et al. 2013).
We conducted the spectral fitting over flaring periods, which are identified as described in Section 3.1. For each source in our sample, we defined a region of interest (ROI) with a radius of 15 degrees centered on the source. The filter expression used is ‘(DATA QUAL¿0)(LAT CONFIG == 1)(angsep(RA source, DEC source, RA_SUN, DEC_SUN)¿15)’, where RA source and DEC source are the right ascension and declination of the source, and RA_SUN and DEC_SUN are those of the Sun. This selection ensures that only good time intervals (GTIs) are considered, and time periods when the source is within 15 degrees of the Sun are excluded to prevent contamination from solar emission.
The -ray sky model was constructed using the 4FGL-DR3 catalog, and the energy range considered was 50 MeV to 300 GeV. We used different PSF event types and zenith angle cuts for different energy ranges to optimize angular resolution and sensitivity:
@@
This strategy provides the best possible angular resolution and minimizes background contamination at low energies, ensuring that our gamma-ray data analysis is as robust as possible. Also, this selection maximizes sensitivity to point sources like blazars, which is crucial for accurate spectral and spatial analyses.
@@
The normalization of the Galactic and isotropic diffuse component was left free. Additionally, the normalization parameter of catalog sources with a test statistic (TS) value greater than 25 within 9 degrees of the ROI center is set free. For sources with TS 500 within 12 degrees, both the normalization and spectral index are allowed to vary, accounting for long-term averaged source properties over the first 12 years of Fermi-LAT data. After an initial fitting, a source-finding step is implemented using gta.find_sources from Fermipy to identify additional unresolved sources in the ROI. Following this, the same TS-based criteria are reapplied to both catalog and newly identified sources during the analysis period, refining the model to account for temporal variability or new source activity. Finally, a global fit is performed to obtain the optimized model parameters.
@@
We then tested the broken power-law (BPL) model as a candidate spectral function. Initially, we fitted a power-law (PL) spectral function and recorded the likelihood of the fit, . Subsequently, we changed the spectral model to a broken power-law(BPL) function, defined as
@@
where is the break energy (in MeV), is the photon index for , and is the photon index for . We iterated the fit by fixing the break energy, , at 100 logarithmic intervals between 0.1 GeV and 10 GeV, calculating the likelihood, , at each step. The photon indices ( and ) are allowed to vary between -5 and 5. At each iteration, we calculated the value of TShardening = 2 ( - ), which quantifies the test statistics of spectral hardening from comparing the BPL model likelihood with the PL model likelihood. Note that this is similar, but more specific, to the TScurv used in the 4FGL catalog for determining whether a curved spectral function is preferred over a PL function (Abdollahi et al. 2020). And while creating the SEDs, we left the parameters of all sources free within 3 degrees. We finally selected the that maximized the TShardening profile.
@@
The choice of a BPL function is both statistically and physically motivated. Flaring periods are characterized by rapid spectral changes and generally short integration times, making smoother functions such as a logparabola less effective for constraining spectral hardening (e.g., Abdo et al. 2010; Ackermann et al. 2011). Furthermore, a BPL, with an abrupt spectral hardening and two independent photon indices, better captures sudden changes in particle acceleration or cooling while providing a clearer quantification of spectral hardening (e.g., Tavecchio et al. 2009).
In summary, we defined spectral hardening as cases where the Fermi-LAT spectrum (50 MeV – 300 GeV) is significantly better fit by a broken power law than by a single power law, with the photon index above the break energy harder than below. This operational definition, consistent with previous Fermi-LAT studies (Abdo et al. 2010; Brown et al. 2017; Rulten et al. 2020; Zeng et al. 2022; Paliya et al. 2025), is applied here as a statistical method to identify localized slope changes, without invoking a specific physical interpretation.